# VesselSAM: Leveraging SAM for Aortic Vessel Segmentation with LoRA and Atrous Attention

### v1

@@ IV-C Quantitative results
In this section, we provide a comprehensive comparison of our proposed method, VesselSAM, against various state-of-the-art (SOTA) models, including UNET[25], UNETR[8], SAM[6], MedSAM[10], SAMMed[9], SAMed[11] and SAM-Adopter[19]. Each method was assessed under the same conditions to ensure a fair comparison, allowing us to accurately evaluate performance metrics such DSC and HD. The results highlight how our approach outperforms the SOTA models in the field, demonstrating its effectiveness in tackling complex medical image segmentation tasks.
@@ IV-C1 Quantitative Evaluation Results for AVT Dataset
The performance metrics for various segmentation methods on the Aortic Vessel Tree (AVT) dataset are presented in Table. I. This comparison encompasses both big and small models, illustrating the effectiveness of each approach across multiple hospitals. VesselSAM demonstrates exceptional segmentation performance, achieving a Dice Similarity Coefficient (DSC) of 93.50% at Dongyang Hospital, 93.25% at Rider Hospital, and 93.02% at Kits Hospital. This performance significantly surpasses that of state-of-the-art methods, including MedSAM and SAMAdp.
The incorporation of Atrous Attention and LoRA mechanisms within VesselSAM has contributed to its high performance, enabling the model to effectively capture multi-scale features essential for precise segmentation in medical imaging. In contrast, the other SAM-based models such as SAMAdp and SAMedAdp struggle with accuracy, as they produce a significant number of false positive regions within their segmentations. This disparity underscores the advantages of VesselSAM in accurately delineating vascular structures amidst challenging imaging contexts, ultimately supporting its utility for clinical applications.
@@ IV-C2 Quantitative Evaluation Results for TBAD Dataset
In comparison, SAM and MedSAM display lower performance, with DSC scores of 79.53% and 92.20%, respectively. Moreover, other models such as SAMAdp and SAMedAdp also exhibit challenges in segmentation accuracy, as evidenced by their lower DSC values. The consistent high performance of VesselSAM across both the AVT and TBAD datasets demonstrates its potential as a valuable tool for medical image segmentation, particularly in complex cases where precision is paramount.
@@ IV-D Qualitative results
To provide a more intuitive comparison, we present the qualitative segmentation results for various models, including VesselSAM, SAM, MedSAM, SAMAdp, SAMedAdp, and SAM-MedIA, as illustrated in the Fig. 3. The top row displays the results for aortic vessel segmentation, while the bottom row highlights the segmentation of true lumen (TL) and false lumen (FL) for Type-B Aortic Dissection (TBAD). In the aortic vessel segmentation task, VesselSAM effectively delineates the vessel structures, capturing intricate details that may be overlooked by other models. The segmentation accurately follows the boundaries of the aorta, demonstrating its robustness in identifying the vessel amidst surrounding tissues. In contrast, SAM struggles with segmentation accuracy, leading to significant misalignments with the ground truth, particularly in the definition of vessel edges. MedSAM demonstrates improved performance compared to SAM but still fails to capture some fine details, resulting in inaccuracies in the vessel’s contour. The models SAMAdp, SAMedAdp, and SAM-MedIA, struggles to accurately capture the true positive vessel areas, resulting in a significant number of false positive regions in their segmentations. While these models provide reasonable outputs, they tend to misidentify surrounding areas as part of the vessel structure.
In the segmentation of TL, FL and FLT in TBAD dataset, VesselSAM continues to excel by accurately capturing the luminal structures. The segmentation closely aligns with the GT, effectively distinguishing the TL, FL and the FLT. For better visualization, only the TL and FL are presented. In contrast, SAM encounters significant challenges, with poor segmentation of the TL, resulting in structural misrepresentations. MedSAM provides an improvement over SAM, but it still exhibits inaccuracies that affect its reliability in clinical applications. Other methods like SAMAdp, SAMedAdp, and SAM-MedIA similarly face challenges in accurately delineating the lumens, with occasional missing segments and imprecise boundaries.
Figure 4: Qualitative visual results on the AVT (Dongyang, Rider, Kits) and TBAD datasets using Bounding Box Prompts. The first column shows the input images, followed by the ground truth (GT) in the second column. Subsequent columns shows the results of the following configurations: ”VesselSAM w/o AMM” refers to the VesselSAM model with the MedSAM backbone but without the Atrous Attention Module (AMM), while ”VesselSAM* w/o AMM” uses the SAM backbone without AMM. ”VesselSAM* with AMM” and ”VesselSAM** with AMM” incorporate AMM with the MedSAM and SAM backbones, respectively.The outputs are color-coded to highlight the true lumen (pink), the GT boundary line (yellow) and the bounding box-prompt (blue).
@@
Dataset VesselSAM* VesselSAM** Atrous Attention Module Dice Score
@@ IV-E Ablation Study
In this ablation study, we aimed to evaluate the effectiveness of different configurations of the VesselSAM (Segment Anything Model) on medical image segmentation tasks, particularly focusing on vessel segmentation. We conducted a series of comprehensive ablation experiments to evaluate the impact of key components in the model, including the Atrous Attention Module and the LoRA rank. These experiments help in understanding how the individual elements contribute to the overall performance of the model in terms of segmentation accuracy, using the Dice score as the primary evaluation metric.
In this ablation study, we aimed to evaluate the effectiveness of different configurations of the Vessel SAM (Segment Anything Model) on medical image segmentation tasks, particularly focusing on vessel segmentation. We conducted a series of comprehensive ablation experiments to evaluate the impact of key components in the model, Firstly compared the performance of two baseline models—Vessel SAM initialized with the MedSAM (medical domain-specific) and SAM (general domain) configurations. Secondly also tested an enhanced model incorporating Low-Rank Adaptation (LoRA) and Atrous Convolution. The objective was to analyze the impact of these variations on segmentation performance, using the Dice score as the primary evaluation metric.
@@ IV-E1 Impact of the Backbone and Atrous Attention Module
In this ablation study, we assess the impact of the backbone architecture and the integration of the Atrous Attention module on the performance of VesselSAM. We compare two configurations: VesselSAM initialized with the MedSAM backbone (VesselSAM*) and the SAM backbone (VesselSAM**), which serve as the baseline models for this analysis. Additionally, we introduce the Atrous Attention module into both configurations to evaluate its effect on segmentation performance.
@@
In comparison, the VesselSAM model with the SAM backbone (VesselSAM** with AAM) also benefits from the Atrous Attention module, but the Dice scores are consistently lower. While these results still reflect an improvement over the baseline model with the Atrous Attention module, they demonstrate that the MedSAM backbone—tailored for medical applications—offers a clear advantage when combined with the Atrous Attention module.
@@ IV-F Limitations and Future Work
In summary, this study shows that domain-specific models (MedSAM) combined with adaptation techniques such as LoRA and Atrous Convolution can outperform general-domain baselines, achieving superior segmentation accuracy. These insights are valuable for optimizing medical image segmentation models, particularly when faced with computational limitations. While VesselSAM demonstrates strong performance in vascular segmentation, there are a few limitations that need to be addressed. One key limitation is the reliance on bounding box prompts, which may not always provide enough detail for complex or ambiguous structures. To improve flexibility and accuracy, future work will incorporate alternative prompt mechanisms, such as text-based prompts, to provide richer, more intuitive guidance for segmentation tasks. Another limitation is the model’s dependence on high-quality input images. Although VesselSAM works well on clean, well-annotated datasets, its performance may decrease with noisy or low-resolution images. Future research will focus on enhancing the model’s robustness by incorporating data augmentation techniques and improving its generalization to diverse real-world imaging conditions.In addition, the integration of visual language models with VesselSAM offers exciting potential for future work. By combining the power of language understanding with visual information, these models could further enhance prompt generation and segmentation accuracy, enabling the model to better handle ambiguous or novel vascular structures with minimal user input. Furthermore, while VesselSAM is currently focused on segmentation of the aortic vessels, its potential for other vascular structures and medical domains has not been fully explored. Expanding the model’s applicability to other areas, such as brain or coronary artery segmentation, will be an important direction for future work.
@@ V Conclusion
In this paper, we introduced VesselSAM, an enhanced version of the Segmentation Anything Model (SAM), designed specifically for aortic vessel segmentation. By incorporating the Atrous Attention Module and Low-Rank Adaptation (LoRA), VesselSAM addresses key limitations of the original SAM, enhancing its ability to capture complex hierarchical features typical of medical images. The Atrous Attention Module enables multi-scale feature extraction, effectively capturing both fine-grained details and broader anatomical context, while LoRA optimizes fine-tuning efficiency by reducing the number of trainable parameters without sacrificing performance.
Extensive experiments on the Aortic Vessel Tree (AVT) and Type-B Aortic Dissection (TBAD) datasets demonstrate that VesselSAM outperforms ViT-based and other SAM-based methods, achieving superior Dice Similarity Coefficient (DSC) and HD scores. The model achieves these results with significantly fewer trainable parameters, reinforcing its role as a parameter-efficient fine-tuning (PEFT) model for medical applications. These findings underline VesselSAM’s superior performance while maintaining computational efficiency, making it particularly valuable for real-world clinical tasks.
VesselSAM offers a robust solution for medical image segmentation, combining high accuracy with computational efficiency. Its ability to generalize across diverse datasets and perform well with minimal additional computational resources makes it highly suitable for clinical applications. Future work will explore further optimizations, including the integration of text-based prompts and visual language models, and expand its use to other medical imaging tasks, ensuring broader applicability in the healthcare domain.

### v2

@@ IV-C Quantitative results
A comprehensive comparison is conducted to evaluate the performance of VesselSAM against various state-of-the-art (SOTA) models, including UNET[31], UNETR[8], SAM[6], MedSAM[10], SAMMed[9], SAMed[11] and SAMAdp[19]. Each method was assessed under identical conditions to ensure a fair comparison, allowing us to accurately evaluate performance metrics such DSC and HD. The results demonstrate that VesselSAM surpasses existing SOTA models, effectively addressing challenges in complex medical image segmentation tasks.
@@ IV-C1 Quantitative Evaluation Results for AVT Dataset
The performance metrics for various segmentation methods on the Aortic Vessel Tree (AVT) datasets, including Dongyang Hospital, Rider Hospital, and KiTs Hospital, are presented in Table. I. This comparison encompasses both big and small models, illustrating the effectiveness of each approach across multiple hospitals. VesselSAM demonstrates exceptional segmentation performance, achieving a DSC of 93.50% at Dongyang Hospital, 93.25% at Rider Hospital, and 93.02% at Kits Hospital. This performance significantly surpasses that of state-of-the-art methods, including MedSAM and SAMAdp.
The incorporation of Atrous Attention module and LoRA mechanisms within VesselSAM has contributed to its high performance, enabling the model to effectively capture multi-scale features essential for precise segmentation in medical imaging. In contrast, models such as SAMMed and SAMed exhibit higher false positive rates, leading to suboptimal segmentation accuracy. This disparity underscores the advantages of VesselSAM in accurately delineating vascular structures amidst challenging imaging contexts, ultimately supporting its utility for clinical applications.
@@ IV-C2 Quantitative Evaluation Results for TBAD Dataset
In comparison, SAM and MedSAM display lower performance, with DSC scores of 79.53% and 92.20%, respectively. Moreover, other models such as SAMMed and SAMAdp also exhibit challenges in segmentation accuracy, as evidenced by their lower DSC values. The consistent high performance of VesselSAM across both the AVT and TBAD datasets demonstrates its potential as a valuable tool for medical image segmentation, particularly in complex cases where precision is paramount.
Figure 3: Qualitative visual results on the AVT (Dongyang, Rider, Kits) and TBAD Datasets under Bounding Box Prompts. The first column shows the input images, followed by the ground truth (GT) in the second column. The subsequent columns show the results of various segmentation models: VesselSAM, SAM, MedSAM, SAMAdp, SAMed, and SAMMed. Each model’s output is overlaid with color-coded regions for true lumen (pink), false lumen (green), and the GT boundary line (yellow). The blue box represents the bounding box-prompt used for segmentation. The images have been zoomed in to enhance visibility.
@@ IV-D Qualitative results
To provide a more intuitive comparison, qualitative segmentation results are presented for VesselSAM and the same models evaluated in the quantitative analysis, as illustrated in Fig. 3. The top row displays the results for aortic vessel segmentation, while the bottom row highlights the segmentation of true lumen (TL) and false lumen (FL) for Type-B Aortic Dissection (TBAD). In the aortic vessel segmentation task, VesselSAM effectively delineates the vessel structures, capturing intricate details that may be overlooked by other models. The segmentation accurately follows the boundaries of the aorta, demonstrating its robustness in identifying the vessel amidst surrounding tissues. In contrast, SAM struggles with segmentation accuracy, leading to significant misalignments with the ground truth, particularly in the definition of vessel edges. MedSAM demonstrates improved performance compared to SAM but still fails to capture some fine details, resulting in inaccuracies in the vessel’s contour. The models SAMMed, SAMed, and SAMAdp, struggles to accurately capture the true positive vessel areas, resulting in a significant number of false positive regions in their segmentations. While these models provide reasonable outputs, they tend to misidentify surrounding areas as part of the vessel structure.
In the segmentation of TL, FL and FLT in TBAD dataset, VesselSAM continues to excel by accurately capturing the luminal structures. The segmentation closely aligns with the GT, effectively distinguishing the TL, FL and the FLT. For better visualization, only the TL and FL are presented. In contrast, SAM encounters significant challenges, with poor segmentation of the TL, resulting in structural misrepresentations. MedSAM provides an improvement over SAM, but it still exhibits inaccuracies that affect its reliability in clinical applications. Other methods like SAMAdp, SAMed, and SAMMed similarly face challenges in accurately delineating the lumens, with occasional missing segments and imprecise boundaries.
Figure 4: Qualitative visual results on the AVT (Dongyang, Rider, Kits) and TBAD datasets using bounding box prompts. The first column shows the input images, followed by the ground truth (GT) in the second column. Subsequent columns show the results of the following configurations: ”VesselSAM* w/o AMM” refers to the VesselSAM model with the MedSAM backbone but without the Atrous Attention Module (AMM), while ”VesselSAM** w/o AMM” employs the SAM backbone without AMM. ”VesselSAM* with AMM” and ”VesselSAM** with AMM” incorporate AMM with the MedSAM and SAM backbones, respectively. The outputs are color-coded to highlight the true lumen (pink), the GT boundary line (yellow), and the bounding box prompt (blue). The images have been zoomed in to enhance visibility.
@@
Dataset VesselSAM* VesselSAM** Atrous Attention Module DSC
@@ IV-E Ablation Study
To evaluate the effectiveness of different configurations of the VesselSAM in medical image segmentation tasks, particularly vessel segmentation, we conducted a series of comprehensive ablation experiments. First, we compared the performance of two baseline models—VesselSAM initialized with the MedSAM (medical domain-specific) and SAM (general domain) configurations. Next, we tested an enhanced model incorporating the Atrous Attention module. The objective was to analyze the impact of these variations on segmentation performance, using the DSC as the primary evaluation metric.
@@ IV-E1 Impact of the Backbone and Atrous Attention Module
To assess the impact of the backbone architecture and the integration of the Atrous Attention module on the performance of VesselSAM. We compare two configurations: VesselSAM initialized with the MedSAM backbone (VesselSAM*) and the SAM backbone (VesselSAM**), which serve as the baseline models for this analysis. Additionally, we introduce the Atrous Attention module into both configurations to evaluate its effect on segmentation performance.
@@
In comparison, the VesselSAM model with the SAM backbone (VesselSAM** with AAM) also benefits from the Atrous Attention module, but the Dice scores are consistently lower. While these results still reflect an improvement over the baseline model with the Atrous Attention module, they demonstrate that the MedSAM backbone tailored for medical applications, offers a clear advantage when combined with the Atrous Attention module.
@@ IV-F Limitations and Future Work
This study demonstrates that domain-specific models, such as MedSAM, can achieve superior segmentation accuracy when enhanced with adaptation techniques like AtrousLoRA and Atrous Attention Module. These findings highlight the importance of parameter efficient fine-tuning strategies for optimizing medical image segmentation, particularly under computational constraints. Despite its strong performance in aortic vessel segmentation, VesselSAM has several limitations that warrant further investigation. One primary limitation is its reliance on bounding box prompts, which may not always provide sufficient contextual information for segmenting complex or ambiguous vascular structures. To improve flexibility and accuracy, future work will explore alternative prompt mechanisms, such as text-based prompts, to offer richer, more intuitive guidance for segmentation tasks.
Another challenge is VesselSAM’s dependency on high-quality input images. While the model performs well on clean, well-annotated datasets, its segmentation accuracy may degrade in noisy, low-resolution, or real-world clinical imaging conditions. To address this, future research will focus on enhancing the model’s robustness through advanced data augmentation techniques and strategies to improve generalization across diverse medical imaging domains.
Furthermore, the integration of visual-language models (VLMs) with VesselSAM presents an exciting direction for future work. By leveraging language-driven prompts, these models could refine segmentation accuracy and enable the system to handle ambiguous or novel vascular structures with minimal user input. Additionally, expanding VesselSAM’s applicability beyond aortic vessel segmentation is crucial. Investigating its performance on other vascular structures, such as coronary arteries, cerebral vessels, and peripheral vasculature, could further enhance its clinical utility across multiple medical domains. By addressing these limitations and exploring these future directions, VesselSAM can evolve into a more generalized, adaptive, and clinically impactful segmentation framework for medical image analysis.
@@ V Conclusion
In this paper, we introduced VesselSAM, an enhanced adaptation of the Segment Anything Model (SAM), specifically designed for aortic vessel segmentation. By integrating AtrousLoRA, a novel combination of Atrous Attention and Low-Rank Adaptation (LoRA), VesselSAM effectively overcomes key limitations of the original SAM, improving its ability to capture complex hierarchical features in medical images. The Atrous Attention Module facilitates multi-scale feature extraction, preserving both fine-grained details and broader anatomical structures, while LoRA optimizes fine-tuning efficiency, significantly reducing trainable parameters without compromising segmentation accuracy.
Extensive evaluations on the Aortic Vessel Tree (AVT) and Type-B Aortic Dissection (TBAD) datasets demonstrate that VesselSAM outperforms state-of-the-art (SOTA) ViT-based and SAM-based models, achieving superior DSC and Hausdorff Distance HD scores. Notably, VesselSAM achieves these results with fewer trainable parameters, reinforcing its position as an efficient Parameter-Efficient Fine-Tuning (PEFT) model for medical imaging applications. These findings highlight its ability to deliver high segmentation accuracy while maintaining computational efficiency, making it highly valuable for real-world clinical deployment.
The VesselSAM offers a robust and scalable solution for vascular image segmentation, demonstrating strong generalization across diverse vascular datasets while maintaining minimal computational overhead. Future work will focus on further enhancing its adaptability, including the integration of text-based prompts and visual language models to enrich segmentation guidance, as well as extending its applicability to other vascular structures and medical imaging tasks. These advancements will further solidify VesselSAM’s role as a versatile and efficient AI-driven tool for clinical and research applications in medical imaging.